Q No 3: a). Altman Corporation has interest expenses of $120,000 annually. Altman’s annual sales are $4 million, its tax rate is 25%, and its net profit margin on sales is 10 percent. What is Altman’s TIE?


b). Back Alley Boys, Inc. had sales of $250,000, cost of goods sold of $80,000, depreciation expense of $27,000, and additions to retained earnings of $33,360. The firm paid out $30,000 in dividends. Assume a 34% income tax rate, what is the times interest earned ratio?


c). Using the following data, compute the value of cost of goods sold for Peterson Brewing:

Current liabilities=$340,000 quick ratio=1.8 inventory turnover=4.0 current ratio=3.3


1
Expert's answer
2021-11-22T02:56:59-0500

a)

NPM = "\frac{Earnings after tax(EAT)} {Sales} = 10\%"


EAT=Sales "\times0.10 = 4000000\times0.10 = \$400000"  


EAT = EBT ( 1 - 0.25) = 0.75 EBT


EBT = "\frac{EAT}{0.75} = \frac{400000}{0.75} =\$ 533333.33"


EBIT = EBT + interest = "533333.33 + 120000 =\$ 653333.33"


TIE = "\frac{EBIT}{ Interest} = \frac{653333.33}{120000} = 5.44"


b)

Earnings before interest and tax (EBIT):

= Sales – COGS – Depreciation

"= \$250,000 – \$80,000 – \$27,000"

"=\$143,000"


Earnings before tax (EBT):

= Additions to retained earnings + dividends


"= \frac{(\$33,360 + \$30,000)}{0.66}"


"=\$96,000"


The sum of the additions in retained earnings and the amount of dividends have been divided by 0.66 to arrive at income before tax (IBT).


The annual Interest expense:

= EBIT – IBT


"= \$143,000 – \$96,000"


"= \$47,000"


Times interest earned ratio (TIE):

"= \frac{\$143,000}{\$96,000}"


"=1.49"


C)

Current assets "= 3.3 \times \$340,000 = \$1,122,000"


Quick ratio = (Current assets - Inventory)"\div" Current liabilities


"1.8 = \frac{(\$1,122,000 - Inventory)}{\$340,000}"


Inventory "= \$510,000"


Inventory Turnover = "\frac{Cost of goods sold }{Inventory}"


Cost of goods sold = "4.0 × \$510,000"


Cost of goods sold = "\$2,040,000"